OpenStudy (anonymous):

Casey and Jeremiah made a simple bike ramp out of a piece of straight plywood that is thirty inches long. If the top of the ramp stands fourteen inches high, what is the length of the ramp along the ground? Round your answer to the nearest inch.

OpenStudy (anonymous):

you use a simple a^2+b^2=c^2 formula. So, 30 in is C and 14 is B. a^2+14^2=30^2